If we know that the roots of the cubic equation x3 - 61x2 -8000 = 0 are in G.P., then find the roots of the equation.

1
Expert's answer
2022-04-25T11:57:46-0400

Consider the substitution: "x=z-a". The equation takes the form: "(z-a)^3-61(z-a)^2-8000=0". We expand the latter and get: "z^3-3az^2+3a^2z-a^3-61(z^2-2az+z^2)-8000=0." Set "a=-\frac{61}{3}" to delete the term that contains "z^2". We get: "z^3+pz+q=0" with "p=-\frac{3721}{3}", "q=-\frac{669962}{27}". Compute the following value "\Delta=\frac{q^2}4+\frac{p^3}{27}\approx83253629.63". One of roots of equation is: "z=(-\frac{q}{2}+\Delta)^{(1/3)}+(-\frac{q}{2}-\Delta)^{(1/3)}\approx42.6813". The corresponding root of the cubic equation is: "x_0=z-a\approx63.0147". Consider the function: "f(x)=x^3-61x^2-8000". The derivative is "f'=3x^2-122x=3x(x-\frac{122}{3})". From the latter we get: "f'>0" for "x\in(-\infty,0)\cup(\frac{122}{3},+\infty)". It means that the function increases on intervals: "(-\infty,0)" and "(\frac{122}{3},+\infty)". "f(0)=-8000". "f(\frac{122}{3})<0". Thus, the function has only one real root. It is on the interval "(\frac{122}{3},+\infty)". It means, that it is possible to present the function as: "f(x)=(x-x_0)((x+\alpha)^2+\beta)"with real "\alpha" and "\beta". The following relations holds: "2\alpha-x_0=-61", "-2\alpha x_0+\alpha^2+\beta=0" We receive: "\alpha=\frac{x_0-61}{2}\approx1.0073", "\beta=2\alpha x_0-\alpha^2\approx125.9398". Other two roots have the form: "x_1=-\alpha-i\sqrt{\beta}\approx-1.0073+11.2223\,i,\quad x_2=-\alpha-i\sqrt{\beta}\approx-1.0073-11.2223\,i."

Thus, roots of equation have the form: "x_0\approx63.0147", "x_1\approx-1.0073+11.2223\,i", "x_2\approx-1.0073-11.2223\,i".